OpenStudy (anonymous):

4^(x-1) = 2^x +8 Any guiding tips would be greatly appreciated!

OpenStudy (anonymous):

This? \[4^{x-1} = 2^{x}+8\]

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

you sure it is not \[4^{x+1}=2^{x+8}\]??

OpenStudy (anonymous):

4x-1+8

OpenStudy (anonymous):

or rather \[4^{x-1}=2^{x+8}\]

OpenStudy (anonymous):

Alex had it right. It's \[4^{x-1} = 2^x + 8\]

OpenStudy (anonymous):

Hmm, honestly I'm not sure how you would solve that...

OpenStudy (zarkon):

\[u=2^x\]

OpenStudy (anonymous):

2^(2x-2) - 2^x - 8 = 0 Or just 2^(2x-2) = 2^x + 2^3

OpenStudy (anonymous):

Not sure from there

OpenStudy (turingtest):

\[\frac14 (2^{2x})-2^x-8=0\]then use the quadratic formula perhaps...?

OpenStudy (zarkon):

yes...\[u=2^x\] ;)

OpenStudy (anonymous):

i dont think so, it's exponential not quadratic.

OpenStudy (anonymous):

Zarkon what do you mean?

OpenStudy (zarkon):

it is quadratic in \(2^x\)

OpenStudy (anonymous):

How? What if x isn't 2?

OpenStudy (turingtest):

\[(2^x)^2-4(2^x)-32=0\]\[2^x=\frac{4\pm\sqrt{4^2+4(32)}}{2}=2\pm6\]

OpenStudy (anonymous):

How'd you do that?

OpenStudy (turingtest):

I listened to Zarkon...

OpenStudy (anonymous):

Well, i didnt know what he was saying in the first place

OpenStudy (turingtest):

I also used your formula 2^(2x-2) - 2^x - 8 = 0 and modified it a little...

OpenStudy (anonymous):

Will you show me?

OpenStudy (turingtest):

\[2^{2x-2} - 2^x - 8 =2^{-2}\cdot2^{2x}-2^x-8\]\[=\frac14(2^x)^2-2^x-8= 0\]multiply by 4 and you get what I wrote.

OpenStudy (anonymous):

Thanks Turing and Zarkon, I got it! It works out once you substitute 2^x and use the guadratic formula. (Don't forget to plug back into 2^x)

OpenStudy (turingtest):

All thanks to Zarkon, I had no idea how to solve it ;)

OpenStudy (anonymous):

Sweet turing, thanks, I still do not understand what zarkon was saying though u = 2^x is u just a random variable?

OpenStudy (zarkon):

random variable...only if x was a random variable ;)

OpenStudy (turingtest):

if you sub 2^x=u you get\[\frac14u^2-u-8=0\]so that sub makes the method clear

OpenStudy (anonymous):

Why couldn't he just say... I thought it was for some formula I'd never heard of, always used a,b,c for quadratic.

OpenStudy (turingtest):

Zarkon is not very chatty, you have to make the most of what he says :D

OpenStudy (anonymous):

I see. Thanks guys

OpenStudy (jamesj):

Yes, Zarkon is famously laconic.

OpenStudy (anonymous):

Perhaps, but helpful for sure! :D